# Revisiting Point Cloud Completion: Are We Ready For The Real-World?

### v2

@@ 6.2 Results
ODGNet 112.6 87.3
ODGNet with 0-dim $\mathbf{PH}$ 103.6 80.1
Table 3: Performance of ODGNet with and without topology regularization.
Figure 8: Left: Corresponding Ground Truth of a partial input to TOPODGnet. Middle: Output generated by TOPODGnet. Right: Output generated by ODGNet. We find that TOPODGnet is able to follow the global topology of the ground truth point cloud better than ODGNet(right), as indicated by the tapering upper half of the output and overall consistency. TOPODGnet shows visible benefits of introducing topological priors.
We demonstrate the effect of topology on the ODGNet backbone. Figure 8 shows the topologically consistent output of TOPODGnet relative to the baseline - which in turn results in a reduced Chamfer distance in Table 3. While both the outputs are still noisy, TOPODGnet(middle) maintains a better topology when matched with the ground truth.
@@ 7 Conclusion
Real-world point clouds have several inherent characteristics that contrast them from synthetic point clouds - non-uniform density, noise, and presence of topological features. Current state-of-the-art methods for point cloud completion do not account for these, thereby failing miserably for real-world point clouds. We introduce the first real world paired point cloud completion dataset- RealPC. It consists of 21 categories across $\sim$ 40,000 pairs. We benchmark several state-of-the-art baselines on RealPC and demonstrate the need to rethink point cloud completion for the real world. We also demonstrate that utilizing Homology-based topological features of real-world point clouds as regularizers can help generate topologically accurate complete point clouds.
@@ 8 Persistent Homology
Figure 9: Left to Right (a to c). Progression of filtration on a point cloud over different spatial resolutions as the distance threshold increases [33]. Creation and destruction of $k-dim$ topological features documented in (d) which is a persistence diagram representing $(b_{i},d_{i})$ pairs - each point corresponds to a $k-dim$ homology that takes birth at $b_{i}$ and dies at $d_{i}$.
@@
These holes, generalized across all dimensions, encapsulate the structural essence of a space. Figure 9 illustrate how homology captures these features, translating the geometry of a space into meaningful topological features [19].
@@
Table 6: Persistence Diagram of nine point clouds (of Table 4 and 5) of the RealPC dataset.
@@
9 Topological Loss Function
The coarse point cloud seeds at the decoder assist the topological module and are used as input for the topological loss. The output of the topological module consists of 0-dim topological features (birth, death) pairs. The loss function in Equation 2 uses these pairs for adjusting the birth and death values of the 0-dim Homology features according to the required skeleton and the available components in the input incomplete point cloud.
$\mathcal{L}_{homo}=-\sum_{i=0}^{n}1\{i>k\}(b_{i}-d_{i})\\ =-\sum_{i=k+1}^{n}(b_{i}-d_{i})$ (2)
The parameter k holds significance here. As explained in the main paper, the input partial point cloud may consist of multiple disconnected components as demonstrated in Figure Figure 6 in the main paper. In such cases, the above loss formulation ensures that the skeleton can have three skeleton components that can (a) cater to the three disconnected components and (b) merge together to form the final complete point cloud skeleton.
Setting the value of k requires manual inspection of the partial complete point clouds. In cases where the majority of points clouds have single components, setting k=1 works fairly well.
This formulation assists in the formation of the 0-dim $\mathcal{PH}$ skeleton easily when the partial point clouds are not continuous (points are in disconnected cluster).
10 Experimental Details
11 Demonstrations
We visually demonstrate nine point clouds from our dataset in Table 4 and 5. These demonstrate the non-uniform sparsity and noise that are natural and intrinsic characteristics of real-world datasets. We demonstrate the Persistence Diagrams of these nine point clouds in Figure 6. For most of the point clouds we observe non-trivial persistence (most points are far from the diagonal). This indicates that real-world point clouds captured in challenging settings exhibit significant zero and one-dimensional topological features which are absent in synthetic point clouds (refer to Table 1 in the main paper). These persistence features are not observed in the persistence diagram of the synthetic datasets as shown in Table 7. Almost all the 0-dimensional topological features (indicated by dot) are along or very close to the diagonal for the synthetic datasets (Table 7) as opposed to RealPC (Table 6).

### v3

@@ 6 Experiments
ODGNet 119 111
TopODGNet 103 80
SnowFlakeNet 60 72
BOSHNet 69 5.4
Table 5: Performance comparison of TopODGNet and BOSHNet against the baseline models.
Figure 9: We find that TopODGNet is able to follow the global topology of the ground truth PC better than ODGNet, as indicated by the tapering upper half of the output and overall consistency. TopODGNet shows visible benefits of introducing topological priors.
Figure 10: BOSHNet, on account of multiple 0-dim $\mathcal{PH}$ priors, is able to reconstruct the incomplete PC reasonably.
We demonstrate the benefit of topological priors on the ODGNet backbone. Figure 9 shows the relative topological consistency of TopODGNet w.r.t. the baseline. This results in a slight reduction of Chamfer distance as shown in Table 5. While both the outputs are still noisy, TopODGNet maintains a better topological consistency w.r.t. the ground truth.
We demonstrate the effectiveness of our Homology Sampler Completion Network, BOSHNet in Figure 10 and Table 5. The Homology Sampler outperforms all baselines by a significant margin for CD-L2. It performs comparable to the best baseline for CD-L1. A vital benefit of our approach is that the network has access to the multiple 0-dim $\mathcal{PH}$ prior backbones from the start of the training, unlike the case with TopODGNet, which depends on the learned sparse seeds to generate the backbone. These seeds are noisy during the start of the training and sharpen only during the later stages of the training.
@@ 7 Conclusion
We discover that real-world PCs in uncontrolled settings have non-uniform density, noise, and rich and versatile topological features (extracted using tools from Algebraic Topology). These are non-existent in existing datasets. Existing methods for PC completion do not account for these and fail miserably for real-world PCs. We introduce a topologically-rich real-world PC completion dataset, RealPC with 21 categories across $\sim$ 40,000 pairs. We benchmark several state-of-the-art baselines on RealPC and demonstrate the need to rethink PC completion for the real-world. We demonstrate that utilizing $\mathcal{PH}$-based topological features as priors for real-world PCs can help in generating topologically accurate complete PCs.
@@ 8 Persistent Homology
Figure 11: (a) to (c) Progression of filtration on a point cloud over different spatial resolutions as the distance threshold increases [36]. (d) Birth and death of $k$-dim topological features documented in the form of a persistence diagram, i.e., $(b_{i},d_{i})$ pairs, so that each point corresponds to a homology which is born at $b_{i}$ and dies at $d_{i}$.
@@
These holes, generalized across all dimensions, encapsulate the structural essence of a space. Figure 11 illustrates how homology captures these features, translating the geometry of a space into meaningful topological features [18].
@@
9 Evaluation Metrics
We evaluate the difference between model reconstructed complete PC and ground truth complete PC using the following metrics.
  • Chamfer Distance (CD): It tries to capture the average mismatch between points in two given point clouds $X,Y\in\mathbb{R}^{3}$ and is given by:
    $d_{CD-L2}(X,Y)=\sum_{x\in X}\min_{y\in Y}||x-y||_{2}^{2}+\sum_{y\in Y}\min_{x\in X}||x-y||_{2}^{2}$
    $d_{CD-L1}(X,Y)=\sum_{x\in X}\min_{y\in Y}|x-y|+\sum_{y\in Y}\min_{x\in X}|x-y|$
  • Hausdorff Distance (HD): If the distance between a point $p$ on a surface $S$ and a surface $S^{\prime}$ is given by:
    $d(p,S^{\prime})=\min_{p^{\prime}\in S^{\prime}}||p-p^{\prime}||_{2}$
    then the HD between $S$ and $S^{\prime}$ is defined as:
    $d_{H}(S,S^{\prime})=\max_{p\in S}d(p,S^{\prime})$
    For the evaluation of our surface reconstruction task (Section 4.3 of the main paper), we use an average of $d_{H}(S,S^{\prime})$ and $d_{H}(S^{\prime},S)$.
10 Scene-Level Datasets
We provide a detailed description of the procedure for creation of RealPC. To create RealPC, as mentioned in Section 4.1 of the main paper, we use four open-source railway datasets: [15] by Hungarian State Railways acquired with a Riegl VMX-450 high-density mobile mapping system; [42] by Wuhan University, in which urban railway dataset was captured using Optech’s Lynx Mobile Mapper System, rural railway dataset with MLS system equipped with HiScan-Z LiDAR sensors, and plateau railway dataset with Rail Mobile Measurement System (rMMS) equipped with a 32-line LiDAR sensor; [1] by SNCF Réseau, the French state-owned railway company; and a catenary arch dataset [49] given by Strukton Rail, captured using Trimble TX8 Terrestrial Laser Scanner (TLS). A few scenes from these datasets have been visualized in Figure 3 in the main paper.
We use four open-source scene-level railway datasets from different countries. These scene-level PCs are annotated into relevant sections such as vegetation, overhead cables, railway tracks, industrial support and transmission structures, ground, etc. For this work, we extract industrial structures from these scene-level PCs. As a result, we obtain a PC with multiple industrial structures as shown in Figure 4 in the main paper - Input to (A). Referring to the same figure, our methodology is divided into five main parts. In (A) we employ HDBSCAN [7], a hierarchical clustering algorithm, to cluster individual industrial structures. HDBSCAN works accurately for noisy and complex data and automatically detects the number of data clusters. In (B), a manual inspection is performed to extract industrial structures which can serve as good ground truth for supervised training. Finally, in (C), (D), and (E), we process these ground truth PCs in three different ways to induce uniform sparsity, non-uniform sparsity, and incompleteness as explained in detail ahead.
For making a partial PC (C), we pick a random viewpoint around a ground truth PC and remove N number of points that are farthest away from this viewpoint. Similarly for sparsifying the same PC non-uniformly (D), we again pick a random viewpoint but this time we assign a probability to each point, which is either proportional or inversely proportional to a point’s cubed distance to this viewpoint. Then, we sample N points based on these probability values. This is repeated for different values of N and for different viewpoints. For uniform sparsification (E), we randomly sample N points and then repeat for different values of N. Simultaneous to these steps, we perform manual noise removal. These three steps are conducted for all ground truth PCs to obtain a paired object-level training dataset.
Steps (B) and (C) have manual interventions. We want to highlight that these interventions make RealPC robust and accurate. Most point clouds extracted from industrial scenes are incomplete due to view occlusions present during acquisition. To ensure that all ground truth point clouds in RealPC are complete in all respects, we manually hand-pick the complete shapes (B). To group samples into different classes, manual inspection and labeling is known to be more accurate as compared to automated processes. We intensively study all available shapes and acquire domain knowledge of the dataset before manually categorizing the point clouds.
RealPC is a real-world object point cloud dataset generated using mostly scene-level parent datasets. Topological properties of RealPC are inherited from the parent datasets. Our proposed method exploits these properties using $\mathcal{PH}$ priors.
A differentiating factor of RealPC is that - unlike existing datasets (which may or may not be simulated) that are captured in extremely controlled environments, RealPC is acquired in uncontrolled real-world multi-sensor industrial settings that have numerous factors of variation, noise and disturbance.
We also highlight in the main paper that several existing datasets (mentioned in L129-132) though not simulated, are captured in extremely controlled environments - unlike real-world settings (e.g. industrial) that have numerous factors of variation, noise and disturbance. These datasets (including KITTI) do not consist of ground truth complete shapes and hence do not consist of partial-complete paired information.
11 Non-Neural Methods
In subsection 4.3 of the main paper, we benchmark our RealPC and ShapeNet against three non-neural network-based tasks: (a) Simplification, (b) Surface reconstruction, and (c) Upsampling. Point cloud simplification reduces point density of a point cloud while preserving its salient features and the overall structure. This enables faster processing, lower storage costs, and efficient analysis for large-scale 3D data. We use Weighted Locally Optimal Projection (WLOP) [31] in the mentioned subsection, an enhancement of the parameterization-free denoising and simplification method known as Locally Optimal Projection (LOP) [34]. While LOP struggles with point clouds that exhibit non-uniform distributions, WLOP addresses this shortcoming by integrating locally adaptive density weights.
Surface reconstruction for point clouds aims to generate a continuous surface from discrete points, enabling the creation of complete 3D models essential for visualization, analysis, and manufacturing. An alpha shape [20] is basically made from a subcomplex of the Delaunay triangulation of a point cloud, with its refinement level ranging from a rough approximation to a highly detailed representation of the point cloud’s surface. For our comparison, we use a range of $\alpha$ values to create a mesh from different point clouds. Mesh visualizations shown in Figure 12 support the average high HD values in the case of RealPC as compared to ShapeNet (Table 2 of the main paper). The meshes formed from RealPC point clouds fail to capture the intricate details of the point cloud at any $\alpha$ value.
Point cloud upsampling increases the density of sparse point clouds, enabling finer surface details and improving the accuracy of reconstruction, analysis, and downstream processing tasks. We adopt an upsampling strategy [32] which uses an edge-aware method to compute normals away from the surface singularities, and then uses the data obtained to resample towards the said singularities with the help of a bilateral projector.
12 Topological Loss
(a) $\alpha=0.05$
(b) $\alpha=0.1$
(c) $\alpha=0.15$
Figure 12: Surface reconstruction of some instances from different categories of ShapeNet and RealPC using alpha shapes.
12.1 Formulation
The coarse point cloud seeds at the decoder assist the topological module and are used as input for the topological loss. The output of the topological module consists of 0-dim topological features (birth, death) pairs. The loss function in Equation 3 below uses these pairs for adjusting the birth and death values of the 0-dim Homology features according to the required skeleton and the available components in the input incomplete point cloud.
The parameter k holds significance here. As explained in the main paper, the input partial point cloud may consist of multiple disconnected components as demonstrated in Figure 6 in the main paper. In such cases, the above loss formulation ensures that the skeleton can have three skeleton components that can (a) cater to the three disconnected components and (b) merge together to form the final complete point cloud skeleton.
Setting the value of k requires manual inspection of the partial point clouds. In cases where the majority of point clouds have single components, setting k=1 works fairly well.
This formulation assists in the formation of the 0-dim $\mathcal{PH}$ skeleton easily when the partial point clouds are not continuous (points are in disconnected clusters).
12.2 PH priors for KITTI scene understanding
We test these topological priors for scene reconstruction as well as completion on the real-world KITTI Odometry dataset (data). We cannot use TopODGNet and its modules (dictionary, seed generation, orthogonality constraint) are optimized for shape completion [6]. For KITTI scenes, we develop a generative model (inspired by DGCNN) and test it with and without 0-dim $\mathcal{PH}$ priors. We use DGCNN [53] because it can easily be used for integrating topological priors as it satisfies the required criterion for integrating $\mathcal{PH}$ priors. It transforms each point of a point cloud using Edge Convolutions into higher dimensions and retains the number of points across each layer. Therefore, sparse seeds point clouds of higher dimensions can easily be extracted from these layers and used for computation of the Vietoris Complex and persistence information. We use Sequence 08 for testing, 03 for validation, and the rest for training DGCNN with the KITTI dataset. The results for the experiments are shown in Table 6.
Without $\mathcal{PH}$ priors With $\mathcal{PH}$ priors
Scene Completion 1.18 1.04
Scene Reconstruction 1.50 1.18
Table 6: Benchmarking result for scene reconstruction and completion using the KITTI dataset using the Chamfer Distance $\downarrow$ metric. All numbers are (×$\times 10^{-3}$).
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 13: Left: Complete GT Middle: Partial Input to Homology Sampler based Model Right: Output. Our Homology Sampler model, on account of multiple 0-dim $\mathcal{PH}$ priors is able to accurately reconstruct the incomplete PC.
13 Experimental Details
For Homology Sampler based network, we train the network for 1000 epochs with Adam optimizer using an initial learning rate of 5e-4. We use a standard Point Cloud Autoencoder backbone based on PointNet.
14 Demonstrations
We visually demonstrate nine point clouds from our dataset in Table 7 and 8. These demonstrate the non-uniform sparsity and noise that are natural and intrinsic characteristics of real-world datasets.
We demonstrate the Persistence Diagrams of these nine point clouds in Figure 9. For most of the point clouds we observe non-trivial persistence (most points are far from the diagonal). This indicates that real-world point clouds captured in challenging settings exhibit significant zero and one-dimensional topological features which are absent in synthetic point clouds (refer to Table 1 in the main paper). These persistence features are not observed in the persistence diagram of the synthetic datasets as shown in Table 10. Almost all the 0-dimensional topological features (indicated by dot) are along or very close to the diagonal for the synthetic datasets (Table 10) as opposed to RealPC (Table 9).
We also show some more examples of BOSHNet completions using RealPC in Figure 13. BOSHNet picks up precise topology of the point clouds and generates decent complete point clouds.
@@ 14 Demonstrations
Table 9: Persistence Diagram of nine point clouds (of Table 7 and 8) of the RealPC dataset.